# Is Long Context All You Need? Leveraging LLM’s Extended Context for NL2SQL Experiments, Analysis and Benchmark Paper

### v4

@@ 1. Introduction
CHASE-SQL1112nd on the leaderboard as of December 17th, 2024. Long context NL2SQL pipeline was used as one of the three candidate generators for its tuned self-consistency (multiple choice and pick) (Pourreza et al., 2024) 74.46
@@
Table 1 compares the execution results accuracy (Ex Acc) of published NL2SQL methods using BIRD benchmark (Li et al., 2024a), the most popular benchmark for NL2SQL testing. Our Long Context pipeline yields very competitive results without any fine-tuning and without generating multiple answer candidates (self-consistency). It is shown that LLMs, especially smaller capacity models, benefit significantly from fine-tuning as it helps the models focus on relevant patterns in specific data domains and SQL generation (Domínguez et al., 2024; Liu et al., 2024b). While fine-tuning has been a dominant approach, in-context learning (ICL) using the latest LLMs is gaining traction as a preferable alternative, matching the performance of fine-tuned models (Nan et al., 2023; Pourreza and Rafiei, 2024). ICL does not require re-training or updating the model parameters and avoids overfitting to specific data domains. This is particularly appealing for a production NL2SQL system where it serves many customers across many specialized domains.
Self-consistency is another popular technique used in state-of-the-art NL2SQL systems (Wang et al., 2023; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024). The idea is to generate multiple output candidates and pick the most likely one according to some rules (e.g., majority voting) or a fine-tuned picker model. This works nicely due to the stochastic nature of LLMs, where different models (and even the same model) generate varying outputs for the same input. Self-consistency is orthogonal to our work, and one can combine the two (Pourreza et al., 2024) for higher accuracy (see Section 6.2 for more discussion). In this work we focus on identifying and providing additional contextual information for ICL, leveraging the extended context window of a long-context LLM.
@@
  • Our experiments show that just adding many relevant examples (by question similarity) selected from any available example pools (e.g., training dataset) does not necessarily improve NL2SQL accuracy significantly, contrary to what we expected from the many-shot-learning paper (Agarwal et al., 2024). Instead, examples that are constructed using similar SQL constructs – as the underlying dev dataset (but without looking at them) – and relevant schema elements from the same target database improve accuracy. With this observation, we generate many such examples synthetically for many-shot ICL for NL2SQL. It is also the first study of many-shot ICL for NL2SQL with hundreds of examples, whereas prior ICL approaches for NL2SQL utilize only a handful (3-5) of demonstrations/examples .
@@ 2. Related Works
One can refine the output through multiple LLM calls, verifying, fixing and rewriting the SQL outputs. It is also common to sample multiple answer candidates and select (self-consistency) the most probable or consistent output to improve the quality (Wang et al., 2023; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024). In this work, we focus on leveraging the long context ICL for a typical NL2SQL pipeline – which is orthogonal and can be used with self-consistency. (Pourreza et al., 2024) uses our long context pipeline to generate candidates along with a fine-tuned picker.
@@ 3.3. Verify
We use the untuned gemini-1.5-pro LLM to verify the correctness of the final output. The model is given the entire database table schema and question (with additional hints, if provided) to judge. We believe that fine-tuning a verifier or a picker for self-consistency(Wang et al., 2023; Pourreza et al., 2024; Zheng et al., 2024) can further improve accuracy. While we focus on leveraging the long context ICL for a typical NL2SQL pipeline with simpler techniques (self-correction, verification without fine-tuning), the aforementioned techniques are orthogonal and can be used in conjunction. We demonstrate how this verification step can impact the final performance and the generation cost in Section 5.
@@ 4.2. Benchmark evaluation
Table 2. Benchmark performance and generation latency $T$ using the full long-context NL2SQL pipeline and with gemini-1.5 models. $T_{g}$ measures per request SQL output generation time, including time spent on self-correction and retries.
@@ 4.4. Many-shot ICL with example selection
Table 4(c)(a) shows that the train examples did not provide much boost and instead resulted in worse performance. This means that the model’s ability to learn from similar examples from BIRD train is quite limited and it can be even distracted by them. Also, notice that the impact of distraction seems to be limited, say going from 1 similar example (11k tokens) to 100 similar examples (50k tokens). To further examine the model’s sensitivity to many irrelevant train examples, we injected 1 ground-truth example from BIRD dev in the “middle” of the example blob (the average context size is omitted as it should be almost identical to the previous table). While the model can select the ground truth example among many other irrelevant ones, the presence of those irrelevant train examples can confuse the model as opposed to just having that ground truth. The recall matters more than precision, but the model is still sensitive to bad examples (low precision) to a limited degree – note that the accuracy with 100 random examples and 1 ground truth is much higher than no examples at all.
It is also important to note that the model is robust to both the position and ordering of the ground-truth example. (Liu et al., 2024a) showed that LLMs tend to emphasize information at the beginning and end of the context window, highlighting the importance of proper ranking and ordering for retrieval-augmented generation (RAG). As in Table 4(c)(b-c), where the GT is injected at different positions within the context window, we observe that with gemini-pro-1.5 the ordering of the examples (and the schema information relocated by the relocation of the examples in (c)) has a negligible impact, and the long-context LLMs can retrieve the relevant ground truth example (also the schema information) from anywhere within its context window.
@@ 5. Ablation Study
Table 10. Further breakdown of mini-ablation study results, Ex Acc (%). As shown in Table 8 disambiguation (w/o synthetic examples) is more helpful for challenging or extra hard questions, while many-shot ICL with synthetic examples (w/o disambiguation) is more helpful for easier questions.
@@ 6.2. Further performance improvement via self-consistency
Self-consistency is another very popular technique for NL2SQL (Wang et al., 2023; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024), where multiple candidates are generated, and the most consistent answer or the best one is selected by a fine-tuned picker. As demonstrated by the recent state-of-the-art approaches (Pourreza et al., 2024; Gao et al., 2024) on the BIRD leaderboard, self-consistency has become a crucial technique for achieving high accuracy in NL2SQL. This involves generating multiple candidates using different generators (a.k.a. NL2SQL pipelines) and each generator trying multiple times, and choosing the best one. One of the state-of-the-arts, CHASE-SQL, in Table 1 uses our long context pipeline as one of the three candidate generators, contributing to its top-tier performance on the BIRD leaderboard. One caveat of self-consistency is that it can quickly become expensive in terms of latency and the number of LLM calls. Self-consistency is orthogonal to our work, and can further improve the NL2SQL performance if used together with our pipeline. Table 11 illustrates the impact of generating multiple output candidates from our long context pipeline using a generation temperature of 0.5. With an oracle picker (i.e., at least one of the candidates is correct per question) we can uplift the accuracy beyond 70%. In (Pourreza et al., 2024; Gao et al., 2024), the candidate sets are generated using multiple pipelines, as different strategies result in more diverse outputs, further boosting the final accuracy of self-consistency.
@@ A.1. Table retrieval simulation
For each example in the benchmark, table retrieval operates as follows: 1) The simulated user view and query interactions narrow the search space down to  1-100 candidates; 2) The candidates are embedded and re-ranked with Gecko (Lee et al., 2024b) for relevance to the user query; 3) A fixed top-k candidates are passed on to the NL2SQL stage. At the time of test, the recall of the first stage was 82%. As such, 82% is the maximum attainable end-to-end recall, and represents a near perfect result from the re-ranking stage.
@@ A.2. Impact of good column selection
TBR Recall (%) 97.69 100
TBR Precision(%) 89.72 100
CR Recall(%) 97.12 100
CR Precision(%) 69.43 100

### v5

@@ Abstract.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/yeounoh/lc_nl2sql.
@@ 1. Introduction
CHASE-SQL1112nd on the leaderboard as of December 17th, 2024. Long context NL2SQL pipeline was used as one of the NL2SQL pipelines to generate multiple output candidates (Pourreza et al., 2024) 74.46
@@
Table 1 compares the execution results accuracy (Ex Acc) of published NL2SQL methods using BIRD benchmark (Li et al., 2024a), the most popular benchmark for NL2SQL testing. Our Long Context pipeline yields very competitive results without any fine-tuning and without generating multiple answer candidates (self-consistency). It is shown that LLMs, especially smaller capacity models, benefit significantly from fine-tuning as it helps the models focus on relevant patterns in specific data domains and SQL generation (Domínguez et al., 2024; Liu et al., 2024b; Zhou et al., 2024). While fine-tuning has been a dominant approach, in-context learning (ICL) using the latest LLMs is gaining traction as a preferable alternative, matching the performance of fine-tuned models (Nan et al., 2023; Pourreza and Rafiei, 2024). ICL does not require re-training or updating the model parameters and avoids overfitting to specific data domains. This is particularly appealing for a production NL2SQL system where it serves many customers across many specialized domains.
Self-consistency is another popular technique used in state-of-the-art NL2SQL systems (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024). The idea is to generate multiple output candidates and pick the most likely one according to some rules (e.g., majority voting) or a fine-tuned picker model. This works nicely due to the stochastic nature of LLMs, where different models (and even the same model) generate varying outputs for the same input. Self-consistency is orthogonal to our work, and one can combine the two (Pourreza et al., 2024) for higher accuracy (see Section 6.2 for more discussion). In this work we focus on identifying and providing additional contextual information for ICL, leveraging the extended context window of a long-context LLM.
@@
  • Our experiments show that simply adding many relevant examples (by question similarity) selected from any available example pools (e.g., training dataset) does not necessarily improve NL2SQL accuracy significantly, contrary to what we expected from the many-shot-learning paper (Agarwal et al., 2024). Instead, examples that are constructed using similar SQL constructs – as the underlying dev dataset (but without looking at them) – and relevant schema elements from the same target database improve accuracy. With this observation, we generate many such examples synthetically for many-shot ICL for NL2SQL. It is also the first study of many-shot ICL for NL2SQL with hundreds of examples, whereas prior ICL approaches for NL2SQL utilize only a handful (3-5) of demonstrations/examples .
@@ 2. Related Works
One can refine the output through multiple LLM calls, verifying, fixing and rewriting the SQL outputs. It is also common to sample multiple answer candidates and select (self-consistency) the most probable or consistent output to improve the quality (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024). In this work, we focus on leveraging the long context ICL for a typical NL2SQL pipeline – which is orthogonal and can be used with self-consistency. (Pourreza et al., 2024) uses our long context pipeline to generate candidates along with a fine-tuned picker.
@@ 3.3. Verify
We use the untuned gemini-1.5-pro LLM to verify the correctness of the final output. The model is given the entire database table schema and question (with additional hints, if provided) to judge. We believe that fine-tuning a verifier or a picker for self-consistency(Wang et al., 2022; Pourreza et al., 2024; Zheng et al., 2024) can further improve accuracy. While we focus on leveraging the long context ICL for a typical NL2SQL pipeline with simpler techniques (self-correction, verification without fine-tuning), the aforementioned techniques are orthogonal and can be used in conjunction. We demonstrate how this verification step can impact the final performance and the generation cost in Section 5.
@@ 4.2. Benchmark evaluation
Table 2. Benchmark performance and generation latency $T$ using the full long-context NL2SQL pipeline and with gemini-1.5 models. $T_{g}$ measures per request output generation time, including self-correction and retries.
@@ 4.4. Many-shot ICL with example selection
Table 4(c)(a) shows that the train examples did not provide much boost and instead resulted in worse performance. This means that the model’s ability to learn from similar examples from BIRD train is quite limited and it can be even distracted by them. Also, notice that the impact of distraction seems to be limited, say going from 1 similar example (11k tokens) to 100 similar examples (50k tokens). To further examine the model’s sensitivity to many irrelevant train examples, we injected 1 ground-truth example from BIRD dev in the “middle” of the example blob (the average context size is omitted as it should be almost identical to the previous table). While the model can select the ground truth example among many other irrelevant ones, the presence of those irrelevant train examples can confuse the model as opposed to just having that ground truth. The recall matters more than precision, but the model is still sensitive to bad examples (low precision) to a limited degree – note that the accuracy with 100 random examples and 1 ground truth is much higher than no examples at all. It is also important to note that the model is robust to both the position and ordering of the ground-truth example. (Liu et al., 2024a) showed that LLMs tend to emphasize information at the beginning and end of the context window, highlighting the importance of proper ranking and ordering for retrieval-augmented generation (RAG). As in Table 4(c)(b-c), where the GT is injected at different positions within the context window, we observe that with gemini-pro-1.5 the ordering of the examples (and the schema information relocated by the relocation of the examples in (c)) has a negligible impact, and the long-context LLMs can retrieve the relevant ground truth example (also the schema information) from anywhere within its context window.
@@ 5. Ablation Study
Table 10. Further breakdown of mini-ablation study results, Ex Acc (%). As shown in Table 8 disambiguation (w/o synthetic examples) is more helpful for challenging or extra hard questions, while many-shot ICL with synthetic examples (w/o disambiguation) is helpful for easier questions.
@@ 6.2. Further performance improvement via self-consistency
Self-consistency is another very popular technique for NL2SQL (Wang et al., 2022; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024), where multiple candidates are generated, and the most consistent answer or the best one is selected by a fine-tuned picker. As demonstrated by the recent state-of-the-art approaches (Pourreza et al., 2024; Gao et al., 2024) on the BIRD leaderboard, self-consistency has become a crucial technique for achieving high accuracy in NL2SQL. This involves generating multiple candidates using different generators (a.k.a. NL2SQL pipelines) and each generator trying multiple times, and choosing the best one. One of the state-of-the-arts, CHASE-SQL, in Table 1 uses our long context pipeline as one of the three candidate generators, contributing to its top-tier performance on the BIRD leaderboard. One caveat of self-consistency is that it can quickly become expensive in terms of latency and the number of LLM calls. Self-consistency is orthogonal to our work, and can further improve the NL2SQL performance if used together with our pipeline. Table 11 illustrates the impact of generating multiple output candidates from our long context pipeline using a generation temperature of 0.5. With an oracle picker (i.e., at least one of the candidates is correct per question) we can uplift the accuracy beyond 70%. In (Pourreza et al., 2024; Gao et al., 2024), the candidate sets are generated using multiple pipelines, as different strategies result in more diverse outputs, further boosting the final accuracy of self-consistency.
@@ 7. Conclusion
Acknowledgements.
We would like to thank David Culler, Hank Levy, Reza Sherkat, Per Jacobsson, Cosmin Arad, Zeke Miller, Eva Sharma, Richard Kuzma, Xiance Si, Zhixian Yan and the anonymous reviewers for their valuable feedback.
@@ A.1. Table retrieval simulation
For each example in the benchmark, table retrieval operates as follows: 1) The simulated user view and query interactions narrow the search space down to  1-100 candidates; 2) The candidates are embedded and re-ranked with Gecko (Lee2024GeckoVT) for relevance to the user query; 3) A fixed top-k candidates are passed on to the NL2SQL stage. At the time of test, the recall of the first stage was 82%. As such, 82% is the maximum attainable end-to-end recall, and represents a near perfect result from the re-ranking stage.
@@ A.2. Impact of good column selection
TBR Recall (%) / Precision(%) 97.69 / 89.72 100 / 100
CR Recall(%) / Precision(%) 97.12 / 69.43 100 / 100